OpenStudy (anonymous):

What is the value of y in the equation 2(3y − 6) = 0? –2 0 2 6

OpenStudy (anonymous):

@gabbyalicorn

OpenStudy (gabbyalicorn):

Ok, do you know what your first step is to solve this problem?

OpenStudy (anonymous):

im sorry but i am absoloutley clueless

OpenStudy (anitas.):

1 Divide both sides by 2 3y−6=0 2 Add 6 to both sides 3y=6 3 Divide both sides by 3 y=63 4 Simplify 63 to 2 so the answer is?

OpenStudy (gabbyalicorn):

No, that's fine. :) The first step is to simplify both sides of the equation. Like so: 2(3y−6)=0 (2)(3y)+(2)(−6)=0(DON'T FORGET TO DISTRIBUTE) 6y+−12=0 6y−12=0

OpenStudy (gabbyalicorn):

x'D why @AnitaS.

OpenStudy (anitas.):

why wut?

OpenStudy (gabbyalicorn):

You stole my medal. I was just playing. CX

OpenStudy (anitas.):

lol Don't worry i gave you one! :DDD

OpenStudy (gabbyalicorn):

oh, ty

OpenStudy (anonymous):

so i think its b right

OpenStudy (anonymous):

oh sorry a

OpenStudy (anonymous):

-2 right

OpenStudy (anitas.):

no none of those

OpenStudy (anitas.):

that u said

OpenStudy (anonymous):

ughhhhhh what isss itttttt can u tell me

OpenStudy (anitas.):

no ur supposed to figure it out so then you will learn.. i gave you simple steps and you should be able to figure it out...

OpenStudy (anitas.):

simplify 63 to 2 and what do you get?

OpenStudy (anonymous):

im between c and d

OpenStudy (anitas.):

ok and what is ur final answer?

OpenStudy (anonymous):

c ??????

OpenStudy (anitas.):

Correct! ^_^

OpenStudy (anonymous):

yeyyeyeyyeyeyeyeyssssss u have no idea how happy i am can u help me with one more

OpenStudy (anitas.):

sure make another post! :)

OpenStudy (anonymous):

thanks i ll tag u